OpenStudy (anonymous):

the total number of eigths in 2^3/4?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

didn't get the question.....

OpenStudy (anonymous):

the total number of eighths in 2^3/4?

OpenStudy (anonymous):

@sirm3d

hartnn (hartnn):

is it (2^3 ) /4 or 2^(3/4) ?

OpenStudy (sirm3d):

i think the problem is "how many eighths in \[\huge \frac{2^3}{4}\]

OpenStudy (anonymous):

yes i guess

hartnn (hartnn):

to find number of eights, u make denominator = 8 and then numerator is your answer

OpenStudy (anonymous):

it can be 2(whole number) 3/4

OpenStudy (anonymous):

so what will be the answer @hartnn

hartnn (hartnn):

if its (2^3)/4 then 8/4 = 16/8 ....so 16 eights if its 2 3/4 then 2 3/4 = 11/4 = 22/8 .....so 22 eights

hartnn (hartnn):

i think its 2ns one.... 22 eights

hartnn (hartnn):

*2nd

OpenStudy (anonymous):

yes